If a spherical conductor of radius 10 cm has potential v at a point distance 5cm from its center, then the potential at point a distance 15 cm from the centre will be

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Haresh
 
The potential inside a conductor is the same everywhere (if it wasn't, potential differences would exist and currents would flow internally until the charges redistributed and made the internal potential the same everywhere). 

So everywhere, including the surface, of the conductor is at a potential of V. 

Outside the sphere, the potential is given by k/r

This means the external potential is inversely proportional to distance from centre. 

15cm is 3/2 times bigger than the radius (10cm), so the potential at 15cm is 3/2 times smaller than the potential at the surface. 

Potential = V/(3/2) = (2/3)V
 
Regards
Arun (AskIITians Forum Expert)